OpenStudy (anonymous):

PLEASEE HELP > GIVIGN MEDALS 8one-fourth pounds of flour are used to make 4one-eighth papier-mâché paste recipes. How much flour is used to make one recipe? >simplest form

OpenStudy (anonymous):

Since 8 times 1/4 = 2, and 4 times 1/8 = 1/2, the problem just means 2 pounds of flour are used to make 1/2 papier mache paste recipe.

OpenStudy (anonymous):

have you read my answer to your message? :)

OpenStudy (anonymous):

yes